# Data Reconstruction: When You See It and When You Don’t

### v1

@@ 1 Introduction
Reasoning about data privacy is crucial in today’s data-driven world. This includes the design of privacy-enhancing technologies aimed at protecting privacy, as well as identifying vulnerabilities of existing methods by conducting “privacy attacks”. The most severe family of attacks is, arguably, reconstruction attacks, where an adversary takes what appears to be benign statistics and reconstructs significant portions of the sensitive data [20]. Such attacks have helped shape the theory of data privacy as well as expose vulnerabilities in existing real-world systems.
@@
This is reminiscent of the current state of affairs in the related concept of differential privacy (DP) [18]: If an algorithm satisfies DP (with small enough privacy parameters), then it is guaranteed to be “safe” in terms of its privacy implications. However, the fact that an algorithm does not satisfy DP does not immediately mean that it is unsafe privacy-wise. Rather, to convince that an algorithm is unsafe, one usually needs to demonstrate an actual attack, such as a reconstruction or a membership attack. In other words, in the context of data privacy, we currently do not have a good definition of what constitutes “privacy”; only definitions of what it means to “protect” privacy, and definitions for what it means to “attack” an algorithm in order to show a privacy breach. We embrace this way of thinking and leverage it towards formally reasoning about reconstruction, as stated in Questions 1 and 1. We elaborate on each of these two questions separately, in Sections 1.1 and 1.2, respectively.
@@ 1.1 Protecting against reconstruction
These considerations led Balle, Cherubin, and Hayes [5] and Cummings et al. [13] to present definitions with the following flavor:
@@ 1.1.1 Towards a new definition
We arrived at Definition 1.1.2 through the lens of data reconstruction. To provide further evidence supporting the paradigm of Narcissus resiliency, we show in Section 2 that Narcissus resiliency captures concepts from cryptography and privacy as special cases. This includes differential privacy, resiliency to membership inference attacks, security of one-way functions, security of encryption schemes, and more. We show that all of these concepts can be stated in the terminology of Narcissus resiliency, where an adversary is “trying to beat itself in its own game”.
@@ 1.2 Identifying reconstruction
However, at the formal level, this definition is meaningless: Because ${\cal A}$ is chosen after the example $x$, then formally, every example $x$ is “extractable” as there exists ${\cal A}$ that (ignores its input) and outputs $x$. Carlini et al. [16] mentioned that in order to prevent such pathological cases, the program ${\cal A}$ should be shorter than $x$ (and therefore ${\cal A}$ cannot “memorize” $x$). But this by itself is not enough. For example, consider the string $x=\underbrace{34\>34\>\ldots\>34}_{1000\text{ times}}$ and the program ${\cal A}$ “Print ‘343434341000100010001000 times”. Namely, this example illustrates that ${\cal A}$ being short is clearly not evidence for not being able to “memorize” $x$. Therefore, the length of ${\cal A}$ cannot be the only criteria for determining its validity, and there should be some connection between the length of the attack ${\cal A}$, and the “complexity” of the target output. Intuitively, if there is no way to encode the target output $x$ using a short program (i.e., to compress it), then by demonstrating a short attacking program that is able to reveal $x$ given $y$, proves that $x$ is really extracted from the model $y$.
@@ 1.2.1 Defining Extraction via Kolmogorov Complexity
What makes the string 343434343434343 less random than 285628563123452? The notion of Kolmogorov complexity (in short, $K$-complexity), introduced by Solomonoff [54], Kolmogorov [37], and Chaitin [11] in the 60s, provides an elegant method for measuring the amount of “randomness” in an individual string. Informally, the $K$-complexity of a string $x$, denoted by $K(x)$, is the length of the shortest program that outputs the string $x$. Intuitively, a string $x$ has high $K$-complexity, if there is no short program that outputs it and halts.
@@
Section 1.2.1 can easily be extended to datasets $S$ (rather than a single example $x$), and in Appendix D we also consider a probabilistic version of the $K_{{\cal L}}$-complexity that allows some error probability, and redefine Section 1.2.1 w.r.t. this version.
@@ 1.2.2 Narcissus-Resiliency prevents non-trivial extraction of training data
1.2.3 Hardness of verifying the validity of reconstruction attacks
The main limitation of Section 1.2.1 is that Item 2 (lower-bounding the Kolmogorov Complexity) is not verifiable, as computing the $K$-complexity is an intractable problem. We could consider a tractable version of it, called time-bounded $K$-complexity, where given a parameter $t$ we consider only programs that halt within $t$ steps ([37, 52, 56, 36]). This, however, would only relax intractability to inefficiency, which does not help in our context where we would like to verify a reconstruction attack efficiently. While in practice the $K$-complexity can be estimated using Heuristics like zlib compression [39], any such Heuristic can fail to determine some highly compressible patterns, and as we demonstrate in Section D.4.1, this problem is indeed inherent assuming that a basic cryptography primitive (pseudo-random generator) exists.
@@ 1.3 Additional related works
Perhaps the most basic privacy violation is memorization. Feldman and Zhang [25], Feldman [23], Brown et al. [3], Brown, Bun, and Smith [4], Livni [40], and Attias et al. [2] theoretically study the necessity of memorization in learning. Roughly speaking, they show that for some tasks, the output of any accurate algorithm must have large mutual information with the training data. A similar result of un and Zhandry [9] show that there exist learning tasks in which memorization is only necessary for efficient learners. While large memorization (i.e., mutual information) implies that the algorithm is not differentially private, it does not imply the existence of an efficient attack. Indeed, except for [9, 2], these works are typically not constructive (i.e., they do not provide an efficient way to translate the memorization into an efficient privacy attack).
@@
Beimel, Nissim, and Omri [8] and Mironov et al. [48] considered a computational relaxation of (the standard, information-theoretic) differential privacy, where they require that the outputs of two executions on neighboring datasets are indistinguishable only from the eyes of an efficient observer. This relaxation turns out to be necessary for fundamental distributed tasks [47, 30] but also for some (artificial) centralized ones [6, 26]. We note that an efficient privacy attack (e.g., membership inference or reconstruction) implies that the algorithm is not computationally differentially private. But in the opposite direction, if an algorithm is not computationally differentially private, the privacy attack, while being efficient, could be very negligible (e.g., the attack might reveal only a single, insignificant, bit of information about a training example, and still violating computational differential privacy).
@@
In membership inference (MI) [55, 59], an adversary is given a target example and its goal is to infer whether it was included in a model’s training set. Most techniques follow a paradigm of measuring some correlation/loss function between the target example and the model and checking if it exceeds some threshold. For example, Dwork et al. [21] showed that any algorithm that given $x_{1},\ldots,x_{n}\in\{-1,1\}^{d}$ outputs $y\in[-1,1]^{d}$ that is sufficiently close to the average $\frac{1}{n}\sum_{i=1}^{n}x_{i}$, is exposed to the MI attack that given a target example $x$, decides ”IN” or ”OUT” based on whether $\langle x,y\rangle$ (the inner-product between $x$ and $y$) exceeds some threshold. The MI literature is very rich, and includes works that perform attacks on fundamental learning tasks (e.g., [53, 21, 51, 1, 2]) along with MI attacks on more complex models (e.g., [55, 59, 33, 10, 46, 51, 57]). We remark that while MI attacks are indeed a reasonable privacy concern in many scenarios, they are weaker than reconstruction attacks because the attacker needs to know the target training example beforehand while without it there might be no way to extract sensitive information.
@@
The results of Dinur and Nissim [20] provided a foundation for rigorously quantifying reconstruction bounds for general query release mechanisms. In their setting, the dataset is binary, and they show that given a few statistical queries, it is possible to reveal $90\%$ of the dataset. Dwork, McSherry, and Talwar [19], Dwork and Yekhanin [22], and Haitner et al. [30] provide improved results in similar settings like [20] but under weaker accuracy assumptions. In more general settings, Guo et al. [27] demonstrated that both Rényi Differential Privacy (DP) and Fisher information offer robust semantic assurances against reconstruction attacks. Balle, Cherubin, and Hayes [5] introduced the concept of reconstruction robustness (ReRo), establishing a link between reconstruction attacks and DP. Hayes, Balle, and Mahloujifar [29] extended [5]’s to analyze reconstruction attacks against DP-SGD. Kaissis et al. [34] furthered this research by examining the connection between the hypothesis testing interpretation of DP (specifically $f$-DP) and reconstruction robustness. From a more practical standpoint, there is a rich literature on reconstruction attacks on various real-world models (e.g., [24, 14, 61, 32, 14, 60, 12, 31]). In Appendix D (recognizing reconstruction), we focus on three of them in order to illustrate the expressiveness of our Section 1.2.1.
@@
Some of the work towards formalizing legal privacy concepts mathematically has taken an approach similar to our “sandwiching” of the concept of reconstructing rather than attempting a precise answer to Question 1. These works did not attempt to model a legal concept exactly but rather define requirements which are clearly stricter or clearly weaker than the those of the legal concept, yet non-trivial so they can substantiate a claim that the use of a certain technology satisfies or does not satisfy the legal requirement. As one example, in their modeling of the FERPA privacy requirement, [^5] FERPA – The Family Educational Rights and Privacy Act (FERPA) – is is a Federal law that protects the privacy in education records. Nissim et al. [49] provided a definition which is stronger than the legal requirement, yet satisfiable by the use of differential privacy, hence providing strong evidence that the use of differential privacy (with appropriate parameters) satisfies the legal requirements. As another example is the modeling of the GDPR requirement of protection against singling out: Cohen and Nissim [15] defined protection against predicate singling out, a concept which is weaker than the legal requirement. Showing that $k$-anonymity does not protect against predicate singling out they hence claimed that it does neither satisfy the legal requirement.
@@ 2 Expressiveness of Narcissus resiliency
In the introduction we presented the definition of Narcissus resiliency (Definition 1.1.2) with the interpretation of “security against reconstruction attacks”, where the function $R$ encapsulates what a “valid reconstruction” means. We now show that with different instantiations of the function $R$, the same framework can be used to capture many other existing security notions. In this section we show this for the notions of membership inference attacks and for predicate singling out. In Appendices A and B we extend this to additional security notions, including differential privacy, security of one-way functions, and security of encryption schemes. The fact that Narcissus resiliency is expressive enough to capture all these (seemingly unrelated) well-established notions enhances our confidence in its application as a security notion against reconstruction attacks.
@@ 2.1 Membership inference as Narcissus resiliency
Membership Inference (MI) attacks [55, 59] are a family of attacks which are applied mostly to machine learning models. The goal of an MI attack is to determine whether a given data record was part of the training data underlying the model or not. More specifically, let ${\cal D}$ be a distribution over data records, and let ${\cal M}$ be an algorithm for analyzing datasets of size $n$. Algorithm ${\cal M}$ is said to be MI-secure if no adversary ${\cal A}$ has a significant advantage in distinguishing between the outputs of the following two experiments:
@@ 2.2 Predicate singling out as Narcissus resiliency
[15] defined a type of privacy attack called Predicate Singling Out (PSO), intended to mathematically formulate the legal concept of “singling out” that appears in the General Data Protection Regulation (GDPR). Concretely, let ${\cal D}$ be a data distribution and let $S\leftarrow{\cal D}^{n}$ be a dataset containing $n$ iid samples from ${\cal D}$. Let ${\cal M}$ be an algorithm that operates on $S$ and return an outcome $y$. A PSO adversary gets $y$ and aims to find a predicate $p$ that “isolates” one record in $S$, meaning that it evaluates to 1 on exactly one record in $S$. Note, however, that without further restrictions this is not necessarily a hard task. In particular, even without looking at $y$, the adversary might choose a predicate $p$ whose expectation over ${\cal D}$ is $1/n$. Such a predicate would isolate a record in $S$ with constant probability. More generally, if the expectation of $p$ is $w$, then the probability that it isolates a record in a fresh dataset of size $n$ is $n\cdot w\cdot(1-w)^{n-1}$. Thus, for the attack to be considered “significant”, the adversary has to succeed in its attack with probability noticeably higher than this. The formal definition is as follows.
@@
Let ${\cal M}:{\cal X}^{n}\rightarrow Y$ be an algorithm. Let $R$ be the relation that takes a dataset $S$ and a predicate $p$, where $R(S,p)=1$ if and only if $p$ evaluates to 1 on exactly one point in $S$. We say that ${\cal M}$ is $(\varepsilon,\delta)$-Narcissus-predicate-singling-out-secure (NPSO-secure) if for every distribution ${\cal D}$ over ${\cal X}$ and for every attacker ${\cal A}$ it holds that
@@ 3.1 Limitation of Section 3
It is important to note that Section 3 is limited to attackers ${\cal B}$ that only see the output $y$ of the mechanism. But some of the real-world attacks also use the training data in order to construct an extraction evidence. E.g., in the Diffusion Model attack of Carlini et al. [12] (see Section D.3), the attacker ${\cal B}$ first processes the input data in order to find captions of “interesting” images, and then define ${\cal A}$ as the short program that simply queries the model on one of these captions. We note that Narcissus-resiliency does not prevent such attacks in general, because ${\cal B}$ uses the training data to generate ${\cal A}$. But Section 3 does imply that it is impossible to achieve non-trivial extraction of training examples, without knowing them in advanced.

### v2

@@ 1 Introduction
[…]
@@
[…]
@@ 1.1 Protecting against reconstruction
[…]
@@ 1.1.1 Towards a new definition
We arrived at Definition 1.1.2 through the lens of data reconstruction. To provide further evidence supporting the paradigm of Narcissus resiliency, we show in Section 2 that Narcissus resiliency captures concepts from cryptography and privacy as special cases. This includes differential privacy, resiliency to membership inference attacks, security of one-way functions, security of encryption schemes, and more. We show that all of these concepts can be stated in the terminology of Narcissus resiliency, where an adversary is “trying to beat itself in its own game”. Furthermore, the paradigm enables us to express strictly weaker protections compared to differential privacy that are still meaningful. To support this claim, in Appendix A we provide an example of a very natural deterministic mechanism (which is clearly not differentially private) that estimates a counting query under Narcissus-resiliency w.r.t. the family of all i.i.d. distributions and a natural reconstruction predicate $R$.
@@ 1.2 Identifying reconstruction
[…]
@@ 1.2.1 Defining Extraction via Kolmogorov Complexity
[…]
@@
Section 1.2.1 can easily be extended to datasets $S$ (rather than a single example $x$), and in Appendix E we also consider a probabilistic version of the $K_{{\cal L}}$-complexity that allows some error probability, and redefine Section 1.2.1 w.r.t. this version.
To provide further evidence supporting Section 1.2.1, we demonstrate in Section E.3 how three different types of real-world attacks [CarliniLargeModels21, HaimVYSI22, CarliniDiffusion23] can be explained using the terminology of Section 1.2.1.
@@ 1.2.2 Narcissus-Resiliency prevents non-trivial extraction of training data
1.2.3 Verifying the validity of reconstruction attacks
The main limitation of Section 1.2.1 is that Item 2 (lower-bounding the Kolmogorov Complexity) is not verifiable, as computing the $K$-complexity is an intractable problem. We could consider a tractable version of it, called time-bounded $K$-complexity, where given a parameter $t$ we consider only programs that halt within $t$ steps ([Kolmogorov68, Sisper83, Trakhtenbrot84, Ko86]). This, however, would only relax intractability to inefficiency, which does not help in our context where we would like to verify a reconstruction attack efficiently. In practice, we can only use Heuristics to gain some level of confidence. E.g., to apply many well-known compression algorithms on $x$ and check that all of them results with a compressed representation that is much longer than $\mathopen{}\mathclose{{}\left|{\cal A}}\right|$. Yet, we remark that any such Heuristic can fail to determine some highly compressible patterns, and as we demonstrate in Section E.4.1, this problem is indeed inherent assuming that a basic cryptography primitive (pseudo-random generator) exists.
@@ 1.3 Additional related works
[…]
@@
[…]
BNO08,MPRV09 considered a computational relaxation of (the standard, information-theoretic) differential privacy, where they require that the outputs of two executions on neighboring datasets are indistinguishable only from the eyes of an efficient observer. This relaxation turns out to be necessary for fundamental distributed tasks [McGregorMPRTV10, HaitnerMST22] but also for some (artificial) centralized ones [BunCV16, GhaziIKKM23]. We note that an efficient privacy attack (e.g., membership inference or reconstruction) implies that the algorithm is not computationally differentially private. But in the opposite direction, if an algorithm is not computationally differentially private, the privacy attack, while being efficient, could be very negligible (e.g., the attack might reveal only a single, insignificant, bit of information about a training example, and still violating computational differential privacy).
@@
[…]
@@
[…]
@@
[…]
@@ 2 Expressiveness of Narcissus resiliency
In the introduction we presented the definition of Narcissus resiliency (Definition 1.1.2) with the interpretation of “security against reconstruction attacks”, where the function $R$ encapsulates what a “valid reconstruction” means. We now show that with different instantiations of the function $R$, the same framework can be used to capture many other existing security notions. In this section we show this for the notions of membership inference attacks and for predicate singling out. In Appendices B and C we extend this to additional security notions, including differential privacy, security of one-way functions, and security of encryption schemes. The fact that Narcissus resiliency is expressive enough to capture all these (seemingly unrelated) well-established notions enhances our confidence in its application as a security notion against reconstruction attacks.
@@ 2.1 Membership inference as Narcissus resiliency
Membership Inference (MI) attacks [shokri2017membership, yeom2018privacy] are a family of attacks which are applied mostly to machine learning models. The goal of an MI attack is to determine whether a given data record was part of the training data underlying the model or not. More specifically, let ${\cal D}$ be a distribution over data records, and let ${\cal M}$ be an algorithm for analyzing datasets of size $n$. Algorithm ${\cal M}$ is said to be MI-secure if no adversary ${\cal A}$ has a significant advantage in distinguishing between the outputs of the following two experiments:
@@ 2.2 Predicate singling out as Narcissus resiliency
[CohenN20] defined a type of privacy attack called Predicate Singling Out (PSO), intended to mathematically formulate the legal concept of “singling out” that appears in the General Data Protection Regulation (GDPR). Concretely, let ${\cal D}$ be a data distribution and let $S\leftarrow{\cal D}^{n}$ be a dataset containing $n$ iid samples from ${\cal D}$. Let ${\cal M}$ be an algorithm that operates on $S$ and return an outcome $y$. A PSO adversary gets $y$ and aims to find a predicate $p$ that “isolates” one record in $S$, meaning that it evaluates to 1 on exactly one record in $S$. Note, however, that without further restrictions this is not necessarily a hard task. In particular, even without looking at $y$, the adversary might choose a predicate $p$ whose expectation over ${\cal D}$ is $1/n$. Such a predicate would isolate a record in $S$ with constant probability. More generally, if the expectation of $p$ is $w$, then the probability that it isolates a record in a fresh dataset of size $n$ is $n\cdot w\cdot(1-w)^{n-1}$. Thus, for the attack to be considered “significant”, the adversary has to succeed in its attack with probability noticeably higher than this. The formal definition is as follows.
@@
Let ${\cal M}:{\cal X}^{n}\rightarrow Y$ be an algorithm. Let $R$ be the relation that takes a dataset $S$ and a predicate $p$, where $R(S,p)=1$ if and only if $p$ evaluates to 1 on exactly one point in $S$. We say that ${\cal M}$ is $(\varepsilon,\delta)$-Narcissus-singling-out-secure if for every distribution ${\cal D}$ over ${\cal X}$ and for every attacker ${\cal A}$ it holds that
@@ 3.1 Limitation of Section 3
[…]